OpenStudy (anonymous):

what is the distance to the nearest tenth, from S(4, -1) to W(-2, 3)

OpenStudy (mokeira):

\[\sqrt{(x _{1}-x_2)^{2}+(y_1-y_2)^{2}}\] \[\sqrt{(4--2)^{2}+(-1-3)^{2}}\] \[\sqrt{(36+16)}\] \[\sqrt{52}=2\sqrt{13}\] Hope you understood

OpenStudy (mokeira):

this is same as 7.2 (to the nearest 10th)

OpenStudy (mokeira):

@cnjbain